# Learning Fair Representations with Kolmogorov-Arnold Networks

### v1

@@ Abstract
In this paper, we try to circumvent these issues by integrating Kolmogorov-Arnold Networks (KANs) within a fair adversarial learning framework. Leveraging the adversarial robustness and interpretability of KANs, our approach enables a balance between fairness and accuracy. To further facilitate this balance, we propose an adaptive penalty update mechanism that dynamically adjusts fairness constraints during the model training. We conduct numerical experiments on two real-world college admissions datasets, across three different optimization strategies. The results demonstrate the efficiency and robustness of KANs by consistently outperforming the baseline fair learning models, and maintaining high predictive accuracy while achieving competitive fairness across sensitive attributes.
@@ Introduction
In recent years, the widespread adoption of Machine Learning (ML) models in high-stakes decision-making domains, such as college admissions, healthcare, and hiring, has underscored the need for ethically aligned, fairness-aware Artificial Intelligence (AI) systems. While modern deep learning (DL) models offer high predictive capacity, they remain vulnerable to amplifying historical bias in real-world datasets (bickel1975sex). In this context, undergraduate college admissions have undergone notable transformations, emphasizing the fostering of fairness and equal opportunity for all applicants. Specific initiatives, such as the anon’s Non-Discriminatory Policy (nondp), which propose to eliminate race and gender considerations, implement $\%$-plans to recognize top-performing students, and remove standardized testing in the admissions process, showcase the system’s efforts at prioritizing inclusivity. Although these approaches represent definite strides towards an equitable system, the socioeconomically marginalized groups continue to face systemic barriers in the admissions process (chetty2023diversifying). The disparity in resource and support available to these groups often result in unequal academic performance. This underscores the critical need for robust algorithmic approaches that ensure a balanced predictive performance while maintaining fairness across protected attributes to prevent disparate outcomes.
@@
To meet the goals of adversarial robustness and model interpretability, we adopt Kolmogorov-Arnold Networks (KANs) (liu2024kan)- a novel architecture grounded in Kolmogorov-Arnold (KA) representation theorem (schmidt2021kolmogorov). Unlike traditional Multi-Layer Perceptron (MLP) models, which employ scalar weights and fixed activation functions, KANs leverage a composition of learnable univariate spline functions. This allows for a flexible function approximation with inherent model interpretability and smoothness. The edge-based spline parameterization offers fine-grained control over feature transformations, making them well-suited for structured tabular domains. In this work, we propose an adversarially trained KAN framework that integrates an adaptive $\lambda$ update policy, aimed at balancing fairness metrics and predictive scores. We validate this framework across two different sets of real-world freshman applications to University of California Irvine (UCI). The numerical experiments and theoretical justification further demonstrate the efficiency of our proposed adversarial learning framework in achieving robust performance under fairness constraints.
@@
In the context of college admissions, several studies, such as (bickel1975sex) and (bhattacharya2017university), depict the need for fairness-aware decision-making. Given the profound societal impact, the admissions process must uphold equitable treatment across socio-demographic groups. Historical bias embedded in the data often results in unfair outcomes, disproportionately impacting the marginalized communities (gandara2024inside). To systematically analyze such disparities, we formalize the admission decision task into a binary classification problem.
In our formulation, the model is trained using $m$ examples $(x_{i},z_{i},y_{i})_{i=1}^{m}$, where each of these is composed of a feature vector $x_{i}\in\mathbb{R}^{n}$, containing $n$ predictors, a binary sensitive attribute $z_{i}$, and a binary label $y_{i}$. These examples are sampled from a training distribution, say, $\Gamma=(X,Z,Y)\sim s$. Let $\mathcal{X}$, and $\mathcal{Y}$ be the feature space and label space respectively. We aim to develop a framework that leverages KANs within an adversarial debiasing setup using an adaptive penalty update mechanism. The goal is to mitigate socioeconomic bias while maintaining predictive performance by effectively balancing the trade-off between fairness and accuracy. We try to approximate the predictor model, $f$, with parameters $\theta_{f}$ in a zero-sum game setup with an adversary, $g$, with parameters $\theta_{g}$. One of the key factors for algorithmic fairness is group fairness. For our case, we are considering two types of fairness definitions: Demographic Parity and $p\%$-Rule. In this section, we present the fairness notions and works that are necessary to ground our proposition.
@@
In standard supervised learning, a neural network (NN) model $f$ is typically trained by minimizing the expected prediction loss, $\mathcal{L}_{\mathcal{Y}}$, without incorporating any fairness constraints. When applied to real-world datasets that reflect historical bias, such models can inadvertently encode and even amplify disparities in predicted outcomes across sensitive subgroups. This violates principles of group fairness, particularly when the sensitive attribute $z$ is correlated with the outcome $y$.
Adversarial debiasing frameworks extend this formulation by introducing an adversary $g_{\theta_{g}}$ that attempts to infer the sensitive attribute $z$ from the predictor’s output $f_{\theta_{f}}(x)$. The predictor $f_{\theta_{f}}$ is simultaneously trained to minimize classification loss while maximizing the adversary’s prediction error, resulting in the following optimization objective:
$\begin{split}\min_{\theta_{f}}\>\mathbb{E}_{(x,z,y)\sim s}\>\mathcal{L}_{\mathcal{Y}}(f_{\theta_{f}}(x),y)]\quad\\ \text{s.t.}\quad\min_{\theta_{g}}\>\mathbb{E}_{(x,z,y)\sim s}\>\mathcal{L}_{\mathcal{Z}}(g_{\theta_{g}}(f_{\theta_{f}}(x)),z)>\epsilon^{\prime}\end{split}$ (4)
where, $\epsilon^{\prime}$ is the fairness threshold. To obtain a better balance between the predictions of the predictor and adversary, (zhang2018mitigating) proposed a more relaxed formulation incorporating a fairness coefficient, $\lambda$:
@@
where $\mathcal{L}_{\mathcal{Z}}$ represents the adversary loss, and $\lambda\in\mathbb{R}^{+}$ controls the degree of fairness for striking a balance between the tradeoff.
While adversarial debiasing has proven effective in reducing group-level disparities, we discuss two persistent challenges that remain. Firstly, the fairness constraints (such as DP) lead to provable loss in accuracy, especially when base rates differ across sensitive groups. According to (zhao2022inherent), any predictor that satisfies exact DP incurs an average classification error that is lower bounded by half the base rate gap between the groups. This formalizes the inherent fairness–accuracy tradeoff that can manifest as instability during model training. Moreover, enforcing strict criteria like DP can make models less robust to data perturbations or distribution shifts. The requirement to equalize output distributions across groups may lead to unreliable models, particularly in high-stake settings.
Secondly, while effective at mitigating group-level bias, this adversarial formulation suffers from a critical limitation: the learned neural predictor (i.e. a DL model) $f_{\theta_{f}}$ remains a black-box function that lacks interpretability. Standard NN do not provide explicit functional representations of feature-wise contributions, making it challenging to analyze decision logic, particularly in sensitive real-world settings such as college admissions.
In contrast, KAN’s spline-based architecture and explicit functional decomposition offer a structured, interpretable, and robust alternative to traditional NN. Moreover, KANs retain the flexibility necessary for learning fair decision boundaries. In this study, we present a theoretical analysis and experimental backing demonstrating the suitability of KAN representations for adversarial learning setup.
@@
Several works, (chetty2023diversifying), (barocas2023fairness), (zimdars2010fairness), (woo2023bias), showcase the critical need for fair learning in the college admissions setup. Although various works like (priyadarshini2023admission), (doleck2020predictive), (waters2014grade), (priyadarshini2025fairfusion) showcase ML implementation in this context, adhering to fairness criteria or prioritizing the socioeconomically marginalized groups in the decision-making is still overlooked. This shortcoming is furthered by the reliance on DL models that are hard to interpret, leading to trust issues (rudin2019stop). In the context of college admissions, interpretability plays a significant role, given the direct impact of these decisions on applicants and their subject to public scrutiny.
Although fair-learning approaches like (zhang2018mitigating), (madras2018learning), (lahoti2020fairness), implement adversarial learning to incorporate fairness constraints, they are defined in the context of DL setups. Even several mathematically grounded works such as, (jia2024aligning) introducing Lipschitz-based fairness criterion, and (chzhen2020fair) that leverage optimal mass transport via Wasserstein barycenter for regression tasks, provide rigorous formulations but are mostly tailored to specific problem settings, limiting their applicability. Secondly, works like (zhao2022inherent), (calders2009building), (kamiran2009classifying) further affirm the inherent fairness-accuracy tradeoffs, cited as one of the major problems in our case study. These underlying challenges motivate our proposed approach. While various works like (kiamari2024gkan), (bodner2024convolutional), (vaca2024kolmogorov), (alter2024robustness) have explored KANs on various domains, to the best of our knowledge, our proposed work is the first to elaborate on KAN-based adversarial debiasing framework.
@@
The Adversarial Debiasing (zhang2018mitigating) framework adopts a zero-sum game approach, as shown in Eq. 5. However, due to the inherent fairness-accuracy tradeoff, as discussed in (zhao2022inherent), improvements in fairness often come at the cost of predictive performance. In particular, (hardt2016equality) shows that enforcing DP can eliminate an otherwise perfect predictor, underscoring the importance of theoretically quantifying this tradeoff in classification tasks. Although several adversarial debiasing techniques based on deep NN models have been proposed, (priyadarshini2024fair), they often suffer from a lack of interpretability. In contrast, we propose utilizing KANs (liu2024kan) as the learning models in an adversarial debiasing framework. Having demonstrated adversarial robustness (alter2024robustness), and structural interpretability, KAN models have a definite edge over MLPs. We further incorporate an adaptive $\lambda$-update policy in the KAN-based adversarial framework with various optimizer settings. This helps enhance the model’s robustness to fairness-accuracy trade-off. Despite the high complexity of the KAN representations, we further affirm their efficacy at generalization without overfitting to fairness constraints. An overview of the proposed adversarial learning framework is presented in Fig. 1.
@@
The KANs (liu2024kan) are inspired by the KA Representation theorem (kolmogorov1957representations), which states that any multivariate continuous function can be written as a finite superposition of continuous univariate functions. A distinctive feature of these networks is their use of learnable univariate spline functions, which replace the scalar weights and fixed activation functions in an MLP. These characteristics improves the model’s capacity for accurate and flexible approximations. A traditional KAN function approximator, $f:[0,1]^{n}\to\mathbb{R}$, can be expressed as:
@@
where the inner and outer univariate functions map as, $\phi_{q,r}:[0,1]\to\mathbb{R}$ and $\Phi_{q}:\mathbb{R}\to\mathbb{R}$ respectively. While the early critiques questioned the practical utility of the KA theorem due to the non-smoothness of inner functions (girosi1989representation), $\phi_{i}$, (liu2024kan) demonstrate its viability in ML by leveraging smoother, sparser compositional structures, extending the network to a deep KAN architecture (lin2017does).
Given, all the functions to be learned are univariate, represented in Eq. 6, each 1D function, $\phi_{i}$, is parametrized as a B-spline curve, say $h(x)$, with learnable coefficients of local B-spline basis functions and can be defined as $h(x)=\sum_{i}c_{i}B_{i}(x)$. Each spline function is a piecewise-polynomial (typically cubic) function, and is $(k+1)$ (or, twice differentiable for cubic splines) times continuously differentiable (girosi1989SPLINE). In the following section, we discuss the structural properties of KANs that make them well-suited for fairness-aware adversarial learning with theoretical justification.
@@
Algorithm 1 Training Algorithm incorporating the Adaptive $\lambda$ Update Policy
Input: Initialize KAN-based classifier $f$ and adversary $g$ with grid $\mathcal{G}$ and order $k$. Training data $(x,y,z)$, classifier learning rate $\eta_{\text{clf}}$, adversary learning rate $\eta_{\text{adv}}$, threshold $\tau$, fairness learning rate $\eta$.
Parameter: Initialize penalty weight $\lambda$
@@
$\mathcal{L}_{\mathcal{Y}}(f(x),y)-\lambda\cdot\mathcal{L}_{\mathcal{Z}}$
@@
14:   Compute gaps: $\delta\leftarrow(\tau-p\%\text{-Rule})/\tau$
15:   Update penalties: $\lambda\leftarrow\text{clip}(\lambda+\eta\cdot\delta,0.1,1.0)$
@@
Theoretical Analysis
Lemma 1.
Given $h_{i}:\mathbb{R}\to\mathbb{R}$ denotes the univariate spline functions used in KAN architecture. Based on our Assumptions 1, each $h_{i}$ is a cubic spline continuous function that is $C^{2}$ smooth (i.e., twice differentiable), and defined over a compact interval, $\mathcal{X}\subset\mathbb{R}$. Hence, by the Mean Value Theorem, there exists a constant $L_{i}>0$ such that:
$|h_{i}(x)-h_{i}(x^{\prime})|\leq L_{i}.|x-x^{\prime}|$
for all $x,x^{\prime}\in\mathbb{R}$. As the overall KAN function, $f$, is a finite summation and composition of these univariate spline functions, it follows that $f$ itself is $L$-Lipschitz-continuous on a bounded input domain (rudin1976principles).
Lemma 2.
Every univariate spline function in KAN is $\beta^{\prime}$-smooth, and the finite sum of spline functions preserves the smoothness. Hence, $f$ is $\beta$-smooth.
Given, $h_{i}$ is twice differentiable implies $|h^{\prime\prime}(x)|\leq\beta^{\prime}$ for all $x\in\mathcal{X}$ for some $\beta^{\prime}\in[0,\infty)$. This proves that $h_{i}$ is $\beta^{\prime}$-smooth. Provided the overall KAN function, $f$, is constructed by combining these univariate spline functions through finite summations and compositions, we note two important properties: firstly, the finite sum of spline functions preserves the smoothness, and secondly, composition of smooth functions also preserves the smoothness (rudin1976principles). Hence, it implies that $f$ is $\beta$-smooth. A full justification of the said properties has been provided in the Appendix.
As shown in Lemma 2, the KAN function, $f$, inherits the smoothness of its constituent univariate functions, and remains $\beta$-smooth on the bounded domain, $\mathcal{X}$. This ensures stable convergence during optimization owing to bounded variations in gradients (schmidt2021kolmogorov). Furthermore, Lemma 1 proves Lipschitz continuity of KAN models, hence ensuring model robustness to small input perturbations (alter2024robustness). These structural properties play a crucial role in stabilizing the adversarial training dynamics and ensure effective debiasing. Further mathematical evidence suggesting the efficiency of KAN structure has been provided in the Appendix.
Adaptive Penalty Mechanism
One of the key challenges in an adversarial debiasing set-up lies in balancing fairness and predictive performance, which is governed by the fairness coefficient $\lambda$. It scales the adversarial loss relative to the predictive loss. However, in a non-adaptive setup, determining $\lambda$ is non-trivial. The process requires extensive manual tuning that is computationally expensive (sattigeri2019fairness), especially given the high network complexity of KANs. To address these limitations, we introduce an adaptive $\lambda$ update mechanism that dynamically adjusts the fairness penalty during training. At each training epoch, we compute the fairness gap and update the penalty weights in proportion to this gap. The mechanism follows a linear rule with a learning rate, $\eta$, and is clipped within a safe range to ensure stability in trade-off as shown in Algorithm 1. To further affirm our claims, we empirically demonstrate the effectiveness of our approach in the following section.
Numerical Experiments
@@
Various works, such as (marcinkowski2020implications) and (alvero2020ai), have documented that socioeconomic factors significantly influence college admission outcomes. To empirically assess the efficacy of our proposed method in mitigating such bias, we conduct experiments on two distinct college admissions dataset collected from freshman applicants to the Dept. of Computer Science, UCI. The first dataset, denoted as $D_{(1)}$, comprises 4,442 application records from the Fall 2021 admission cycle, and the second dataset, denoted as $D_{(2)}$, includes 26,243 application records spanning admission cycles from Fall 2018 to Fall 2024. Each of the datasets contains approximately 140 features, encompassing demographic, academic records, high school information, and essay question responses (priyadarshini2023admission). The broad spectrum of features facilitates holistic evaluation, a comprehensive approach to college admissions that ensures applicants are assessed based on a wide range of criteria. For the study, after a systematic analysis of input features, we consider two binary sensitive attributes: low-income status and first-generation flag, which directly affect the socioeconomic demographics. We justify this choice by empirically demonstrating the presence of bias in model predictions using Kernel Density Estimation (KDE) plots, which are included in the Appendix. Furthermore, we use the final read score as the target variable for the binary classification task. To quantitatively evaluate model fairness, we utilize two fairness metrics, $p\%$-Rule, and DP. In addition, to assess the predictive performance of our model, we employ Accuracy and AUROC scores. Next, we incorporate three different types of optimizers, namely Adam (kingma2014adam), Optimistic Adam (OAdam) (daskalakis2017training), and ADOPT (taniguchi2024adopt), for each training configuration. Despite KAN’s structural robustness and theoretical guarantees in adversarial settings, we empirically explore different optimization strategies to enhance model convergence and stability during the debiasing process.
@@
  • Baseline $B_{(1)}$ is a standard adversarial debiasing framework using a fully connected feedforward neural network (FFNN) as the classifier and adversary models,
  • Baseline $B_{(2)}$ is a state-of-the-art implementation of exponential gradient-based debiasing model (agarwal2018reductions),
  • Baseline $B_{(3)}$ is a state-of-the-art method based on the ROAD (Robust Optimization for Adversarial Debiasing) framework (grari2023fairnessROAD).
@@
In this section, we provide a comprehensive comparison of our proposed KAN-based adversarial debiasing framework with an adaptive $\lambda$ update mechanism across two datasets, $D_{(1)}$ and $D_{(2)}$, using three distinct optimizer settings. In the experimental setup, as showcased in Table 1, we use a fixed B-spline order of $k=3$. This choice is consistent with (liu2024kan), that highlight it as striking an effective balance between expressiveness, smoothness, and training stability. Table 1(a) reports the key performance metrics, including classification accuracy, AUROC, $p\%$-Rule, and DP for the two binary sensitive attributes, and Table 1(b) lists the necessary notations.

### v3

@@ Abstract
To circumvent these issues, we propose integrating Kolmogorov-Arnold Networks (KANs) within a fair adversarial learning framework. Leveraging the adversarial robustness and interpretability of KANs, our approach facilitates stable adversarial learning. We derive theoretical insights into the spline-based KAN architecture that ensure stability during adversarial optimization. Additionally, an adaptive fairness penalty update mechanism is proposed to strike a balance between fairness and accuracy. We back these findings with empirical evidence on two real-world admissions datasets, demonstrating the proposed framework’s efficiency in achieving fairness across sensitive attributes while preserving predictive performance.
@@ Introduction
In recent years, the widespread adoption of Machine Learning (ML) models in high-stakes decision-making domains, such as college admissions, healthcare, and hiring, has underscored the need for ethically aligned, fairness-aware Artificial Intelligence (AI) systems. While modern deep learning (DL) models offer high predictive capacity, they remain vulnerable to amplifying historical bias in real-world datasets (bickel1975sex). In this context, undergraduate college admissions have undergone notable transformations, emphasizing the need for fairness and equal opportunity for all applicants. Recent shifts in admission policies, such as the University of California’s Non-Discriminatory Policy (nondp), strive towards non-discriminatory review process with elimination of standardized testing. Although these approaches represent definite strides towards an equitable system, the socioeconomically marginalized groups continue to face systemic barriers in the admissions process (chetty2023diversifying). The disparity in resources and support available to these groups often result in unequal academic performance. This underscores the critical need for robust algorithmic approaches that ensure a balanced predictive performance while maintaining fairness across protected attributes to prevent disparate outcomes.
@@
To meet the goals of adversarial robustness and model interpretability, we adopt Kolmogorov-Arnold Networks (KANs) (liu2024kan) - a novel architecture grounded in Kolmogorov-Arnold (KA) representation theorem (schmidt2021kolmogorov). Unlike traditional Multi-Layer Perceptron (MLP) models, which employ scalar weights and fixed activation functions, KANs leverage a composition of learnable univariate spline functions. This allows for a flexible function approximation with inherent model interpretability and smoothness. The edge-based spline parameterization offers fine-grained control over feature transformations, making them well-suited for structured tabular domains. In this work, we propose an adversarially trained KAN framework with an adaptive $\lambda$ update policy, aimed at balancing fairness metrics and predictive scores. We validate this framework across two different sets of real-world freshman applicants data to the University of California Irvine (UCI). Theoretical analysis and numerical experiments further demonstrate the efficiency of our proposed framework in achieving robust performance under fairness constraints while outperforming state-of-the-art (SOTA) baseline models.
@@
In the context of college admissions, several studies, (bickel1975sex), (bhattacharya2017university), depict the need for fairness-aware decision-making. Given the profound societal impact, the admissions process must uphold equitable treatment across socio-demographic groups. Historical bias embedded in the data often results in unfair outcomes, disproportionately impacting the marginalized communities (gandara2024inside). To systematically analyze such disparities, we formalize the admission decision task into a binary classification problem.
In our formulation, the model is trained using $m$ examples $(x_{i},z_{i},y_{i})_{i=1}^{m}$, where each of these is composed of a feature vector $x_{i}\in\mathbb{R}^{n}$, containing $n$ predictors, a binary sensitive attribute $z_{i}$, and a binary label $y_{i}$. These examples are sampled from a training distribution, say, $\Gamma=(X,Z,Y)\sim s$. Let $\mathcal{X}$, and $\mathcal{Y}$ be the feature space and label space respectively. We aim to develop a framework that leverages KANs within an adversarial debiasing setup using an adaptive penalty update mechanism. The goal is to mitigate socioeconomic bias while maintaining predictive performance by effectively balancing the trade-off between fairness and accuracy. We try to approximate the predictor model, $f$, with parameters $\theta_{f}$ in a zero-sum game setup with an adversary, $g$, with parameters $\theta_{g}$. One of the key factors for algorithmic fairness is group fairness. For our case, we consider two types of fairness definitions: Demographic Parity and $p\%$-Rule. In this section, we present the fairness notions and works that are necessary to ground our proposition.
@@
In standard supervised learning, a neural network (NN) model, $f$, is typically trained by minimizing the expected prediction loss, $\mathcal{L}_{\mathcal{Y}}$, without incorporating any fairness constraints. When applied to real-world datasets that reflect historical bias, such models can inadvertently encode and even amplify disparities in predicted outcomes across sensitive groups. This violates principles of group fairness, particularly when the sensitive attribute, $z$, is correlated with the outcome, $y$.
Adversarial debiasing framework extends this formulation by introducing an adversary $g_{\theta_{g}}$ that attempts to infer the sensitive attribute $z$ from the predictor’s output $f_{\theta_{f}}(x)$. The predictor $f_{\theta_{f}}$ is simultaneously trained to minimize classification loss while maximizing the adversary’s prediction error, resulting in the following optimization objective:
$\begin{split}\min_{\theta_{f}}\>\mathbb{E}_{(x,z,y)\sim s}\>\mathcal{L}_{\mathcal{Y}}(f_{\theta_{f}}(x),y)\quad\\ \text{s.t.}\quad\min_{\theta_{g}}\>\mathbb{E}_{(x,z,y)\sim s}\>\mathcal{L}_{\mathcal{Z}}(g_{\theta_{g}}(f_{\theta_{f}}(x)),z)>\epsilon^{\prime}\end{split}$ (4)
where, $\epsilon^{\prime}$ denotes the fairness threshold, and $\mathcal{L}_{\mathcal{Z}}$ represents the adversary loss. To obtain a better balance between the predictions of the predictor and adversary, (zhang2018mitigating) proposes a more relaxed formulation incorporating a fairness coefficient, $\lambda$:
@@
where, $\lambda\in\mathbb{R}^{+}$ controls the degree of fairness for striking a balance between the tradeoff.
While adversarial debiasing has proven effective in reducing group-level disparities, we discuss two persistent challenges that remain. Firstly, fairness constraints (such as DP) lead to provable loss in accuracy, especially when base rates differ across sensitive groups. According to (zhao2022inherent), any predictor that satisfies exact DP incurs an average classification error that is lower bounded by half the base rate gap between the groups. This formalizes the inherent fairness–accuracy tradeoff that can manifest as instability during model training. Moreover, enforcing strict criteria like DP can make models less robust to data perturbations or distribution shifts. The requirement to equalize output distributions across groups may lead to unreliable models, particularly in high-stake settings. Secondly, while effective at mitigating group-level bias, the adversarial formulation suffers from a critical limitation: the learned neural predictor $f_{\theta_{f}}$ remains a black-box function that lacks interpretability. Standard NNs do not provide explicit functional representations of feature-wise contributions, making it challenging to analyze decision logic, particularly in sensitive real-world settings such as college admissions.
In contrast, KAN’s spline-based architecture and explicit functional decomposition offer a structured, interpretable, and robust alternative to traditional NN. Moreover, KAN retains the flexibility necessary for learning fair decision boundaries. In this study, we present a theoretical analysis and experimental backing demonstrating the suitability of KAN representations for adversarial learning setup.
@@
Several works, (chetty2023diversifying), (barocas2023fairness), (zimdars2010fairness), (woo2023bias), showcase the critical need for fair learning in the college admissions setup. Although various works like (doleck2020predictive), (waters2014grade) demonstrate ML implementation in this context, adhering to fairness criteria or prioritizing the socioeconomically marginalized groups in the decision-making is still overlooked. This shortcoming is furthered by the reliance on DL models that are hard to interpret, leading to trust issues (rudin2019stop). In the context of college admissions, interpretability plays a significant role, given the direct impact of these decisions on applicants and their subject to public scrutiny.
Although fair-learning approaches like (zhang2018mitigating), (madras2018learning), (lahoti2020fairness), (priyadarshini2025fairfusion), implement adversarial learning to incorporate fairness constraints, they are defined in the context of DL setups. Even several mathematically grounded works such as, (jia2024aligning) introducing Lipschitz-based fairness criterion, and (chzhen2020fair) that leverage optimal mass transport for regression tasks, provide rigorous formulations but are mostly tailored to specific problem settings, limiting their applicability. Secondly, works like (zhao2022inherent), (calders2009building), (kamiran2009classifying) further affirm the inherent fairness-accuracy tradeoffs, cited as one of the major problems in our case study. These underlying challenges motivate our proposed approach. While various works like (kiamari2024gkan), (bodner2024convolutional), (vaca2024kolmogorov), (alter2024robustness) have explored KANs on various domains, to the best of our knowledge, our proposed work is the first to elaborate on KAN-based adversarial debiasing framework.
@@
The Adversarial Debiasing (zhang2018mitigating) framework adopts a zero-sum game approach, as shown in Eq. 5. However, due to the inherent fairness-accuracy tradeoff, as discussed in (zhao2022inherent), improvements in fairness often comes at the cost of predictive performance. In particular, (hardt2016equality) shows that enforcing DP can eliminate an otherwise perfect predictor, underscoring the importance of theoretically quantifying this tradeoff in classification tasks. Although several adversarial debiasing techniques based on deep NN models (priyadarshini2024fair) have been proposed they often suffer from lack of interpretability. In contrast, we propose utilizing KAN as the learning model in an adversarial debiasing framework. Having demonstrated adversarial robustness (alter2024robustness), and structural interpretability, KAN models have a definite edge over MLPs. We further incorporate an adaptive $\lambda$-update policy in the KAN-based adversarial framework with various optimizer settings. This helps enhance the model’s robustness to fairness-accuracy trade-off. Despite the high complexity of the KAN representations, we further affirm their efficacy at generalization without overfitting to fairness constraints. An overview of the proposed adversarial learning framework is presented in Figure 1.
@@
The KANs (liu2024kan) are inspired by the KA Representation theorem (kolmogorov1957representations), which states that any multivariate continuous function can be written as a finite superposition of continuous univariate functions. A distinctive feature of these networks is their use of learnable univariate spline functions in place of the scalar weights and fixed activation functions of traditional MLPs. These characteristics improve the model’s capacity for accurate and flexible approximations. A traditional KAN function approximator, $f:[0,1]^{n}\to\mathbb{R}$, can be expressed as:
@@
where the inner and outer univariate functions map as, $\phi_{q,r}:[0,1]\to\mathbb{R}$ and $\Phi_{q}:\mathbb{R}\to\mathbb{R}$ respectively. While the early critiques questioned the practical utility of the KA theorem due to the non-smoothness of inner functions (girosi1989representation), (liu2024kan) demonstrate its viability in ML by leveraging smoother, sparser compositional structures, extending the network to a deep KAN architecture (lin2017does).
Given, all the functions to be learned are univariate, represented in Eq. 6, each 1D function, $\phi_{i}$, is parametrized as a B-spline curve, say $h(x)$. It can be defined as $h(x)=\sum_{i}c_{i}B_{i}(x)$, where $c_{i}$s are the learnable coefficients of the local B-spline basis functions. Each spline function is a piecewise-polynomial (typically cubic) function, and is $(k+1)$ (or, twice differentiable for cubic splines) times continuously differentiable (girosi1989SPLINE). In the following section, we discuss the structural properties of KANs that make them well-suited for fairness-aware adversarial learning with theoretical justifications.
@@
Theoretical Analysis
Lemma 1.
Proof.
Given $h_{i}:\mathbb{R}\to\mathbb{R}$ denotes the univariate spline functions used in KAN architecture. Based on our Assumption 1, each $h_{i}$ is a cubic spline continuous function that is $C^{2}$ smooth (i.e., twice differentiable), and defined over a compact interval ($\subset\mathbb{R}$). Hence, by the Mean Value Theorem, there exists a constant $L_{i}>0$ such that:
$|h_{i}(x)-h_{i}(x^{\prime})|\leq L_{i}.|x-x^{\prime}|$
for all $x,x^{\prime}\in\mathbb{R}$. As the overall KAN function, $f$, is a finite summation and composition of these univariate spline functions, it follows that $f$ itself is $L$-Lipschitz continuous on a bounded input domain (rudin1976principles). ∎
Algorithm 1 Adversarial training with adaptive $\lambda$ policy
Input: Initialize KAN-based classifier $f$ and adversary $g$ with grid $\mathcal{G}$ and order $k$. Training data $(x,z,y)$, classifier learning rate $\eta_{\text{clf}}$, adversary learning rate $\eta_{\text{adv}}$, threshold $\tau$, fairness learning rate $\eta$.
Parameter: Initialize fairness penalty weight $\lambda$
@@
$\mathcal{L}_{\mathcal{Y}}(f(x),y)-\lambda\cdot\mathcal{L}_{\mathcal{Z}}(g(f(x)),z)$
@@
14:   Compute gap: $\delta\leftarrow(\tau-p\%\text{-Rule})/\tau$
15:   Update penalty: $\lambda\leftarrow\text{clip}(\lambda+\eta\cdot\delta,0.1,1.0)$
@@
Lemma 2.
Every univariate spline function, in KAN is $\beta_{i}$-smooth, and the finite sum of spline functions preserves the smoothness. Hence, $f$ is $\beta$-smooth.
Proof.
Given, $h_{i}$ is twice differentiable implies $|h_{i}^{\prime\prime}(x)|\leq\beta_{i}$ for all $x\in\mathcal{X}$, for some $\beta_{i}\in[0,\infty)$. This proves that $h_{i}$ is $\beta_{i}$-smooth. Provided the overall KAN function, $f$, is constructed by combining these univariate spline functions through finite summations and compositions, we note two important properties: firstly, the finite sum of spline functions preserves smoothness, and secondly, the composition of smooth functions preserves smoothness (rudin1976principles). Hence, it implies that $f$ is $\beta$-smooth. A full justification of the said properties has been provided in the Appendix. ∎
As shown in Lemma 2, the KAN function, $f$, inherits the smoothness of its constituent univariate functions, and remains $\beta$-smooth on the bounded domain, $\mathcal{X}$. This ensures stable convergence during optimization owing to bounded variations in gradients (schmidt2021kolmogorov). Furthermore, Lemma 1 proves Lipschitz continuity of KAN models, hence ensuring model robustness to small input perturbations (alter2024robustness). These structural properties play a crucial role in stabilizing the adversarial training dynamics and ensure effective debiasing. Further mathematical evidence suggesting the efficiency of KAN structure in an adversarial debiasing setup has been provided in the Appendix.
Adaptive Penalty Mechanism
One of the key challenges in an adversarial debiasing setup lies in balancing fairness and predictive performance, which is governed by the fairness coefficient $\lambda$. It scales the adversarial loss relative to the predictive loss. However, in a non-adaptive setup, determining $\lambda$ is non-trivial. The process requires extensive manual tuning that is computationally expensive (sattigeri2019fairness), especially given the high network complexity of KANs. To address these limitations, we introduce an adaptive $\lambda$ update mechanism that dynamically adjusts the fairness penalty during training. At each training epoch, we compute the fairness gap and update the penalty weights in proportion to this gap. The mechanism follows a linear rule with a learning rate, $\eta$, and is clipped within a safe range to ensure stability in trade-off as shown in Algorithm 1. To further affirm our claims, we empirically demonstrate the effectiveness of our approach in the following section.
Numerical Experiments
@@
Various works, (marcinkowski2020implications), (alvero2020ai), have documented socioeconomic factors to significantly influence college admission outcomes. To empirically assess the efficacy of our proposed method in mitigating such bias, we conduct experiments on two distinct college admissions dataset collected from freshman applicants to the Dept. of Computer Science, UCI. The first dataset, denoted as $D_{(1)}$, comprises of 4,442 application records from the Fall 2021 admission cycle, and the second dataset, denoted as $D_{(2)}$, includes 26,243 application records spanning admission cycles from Fall 2018 to Fall 2024. Each of the datasets contain approximately 140 features, encompassing demographic, academic records, high school information, and essay question responses (priyadarshini2023admission). The broad spectrum of features facilitates holistic evaluation, a comprehensive approach to college admissions that ensures applicants are assessed based on a wide range of criteria. For the study, after a systematic analysis of input features, we consider two binary sensitive attributes: low-income status and first-generation flag, which directly affect the socioeconomic demographics. We justify this choice by empirically demonstrating the presence of bias in model predictions using Kernel Density Estimation (KDE) plots, which are included in the Appendix. Furthermore, we use the final read score as the target variable for the binary classification task. To quantitatively evaluate model fairness, we utilize two fairness metrics, $p\%$-Rule, and DP. In addition, to assess the predictive performance of our model, we employ Accuracy and AUROC scores. Next, we incorporate three different types of optimizers, namely Adam (kingma2014adam), Optimistic Adam (OAdam) (daskalakis2017training), and ADOPT (taniguchi2024adopt), for each training configuration. Despite KAN’s structural robustness and theoretical guarantees in adversarial setting, we empirically explore different optimization strategies to enhance model convergence and stability during the debiasing process. Further discussions have been included in the Appendix.
@@
  • Baseline $B_{(1)}$ is a standard adversarial debiasing framework using a fully connected feedforward neural network (FFNN) as the classifier and adversary models (priyadarshini2024fair),
  • Baseline $B_{(2)}$ is a SOTA implementation of exponential gradient-based debiasing model (agarwal2018reductions),
  • Baseline $B_{(3)}$ is a SOTA method based on the ROAD (Robust Optimization for Adversarial Debiasing) framework (grari2023fairnessROAD).
@@
In this section, we provide a comprehensive comparison of our proposed KAN-based adversarial debiasing framework with an adaptive $\lambda$ update mechanism across the two datasets, $D_{(1)}$ and $D_{(2)}$, using three distinct optimizer settings. In the experimental setup, as showcased in Table 1, we use a fixed B-spline order of $k=3$. This spline order choice is consistent with (liu2024kan), that highlight it as striking an effective balance between expressiveness, smoothness, and training stability. Table 1(a) reports the key performance metrics, including classification accuracy, AUROC, $p\%$-Rule, and DP across the two binary sensitive attributes, and Table 1(b) lists the necessary notations.